# Generative Photography: Scene-Consistent Camera Control for Realistic Text-to-Image Synthesis

### v2

@@ 4.3 Ablation study
Contrastive Camera Encoder. To assess the impact of contrastive network design, we removed inter-frame differences in the contrastive camera encoder. Table 2 (w/o contrastive) suggests that incorporating different information enhances the model’s accuracy and consistency.
@@ 5 Conclusions
Generative photography opens the door to a new paradigm in photography where the generated contents are consistent with the intrinsic camera physics. Our solution is based on two concepts: dimensionality lifting and contrastive camera learning. It generates physically realistic camera effects while maintaining a high degree of scene consistency. Generative photography has the potential to reduce the post-processing burden in human-operated photography fundamentally. In addition, it offers a fresh perspective for generative models to understand the world better.
Supplementary Materials
@@ 6 Introduction
This supplementary material provides additional discussions and details on the construction of contrastive data (Section 7), network design (Section 8), evaluation metrics (Section 9), and more visual results (Section 10).
To better illustrate the continuity and effects of camera intrinsic setting control, we highly recommend readers view the Videos/GIFs provided in the project page at https://generative-photography.github.io/project/.
7 More Details of Building Contrastive Data
Our contrastive data pipeline dynamically generates training data by storing only base images and scene descriptions. Camera settings are sampled during training and simulated on-the-fly using physical principles, producing contrastive multi-frame data without pre-storing large video files. This also ensures continuous sampling of training data.
We provide below additional key considerations for constructing contrastive datasets for each type of camera setting, along with sample demonstrations.
7.1 Contrastive Data for Bokeh Rendering
@@
7.2 Contrastive Data for Focal Length
In the real world, obtaining a set of images of the same scene at multiple focal lengths is highly cumbersome, with a lack of perfect alignment between the images, and the achievable focal length range is limited [75]. In this paper, we reference the level-of-detail [63, 40] approach and compute the field-of-view (FoV) ratio of the desired focal length relative to the base image focal length. This ratio is then used for center cropping to approximate the actual continuous optical zoom process. In this subsection, we compare the performance of our method with that of actual optical zoom. A camera’s field-of-view (FoV) can be expressed in terms of the focal length $f$ and the sensor dimensions (typically width $w$ or height $h$). The formulas are as follows: Horizontal FoV:
@@ 7.2 Contrastive Data for Focal Length
7.3 Contrastive Data for Shutter Speed
@@
7.4 Contrastive Data for Color Temperature
@@
8 More Details of Contrastive Camera Encoder
In the Contrastive Camera Encoder, an important aspect is the incorporation of the differences in camera setting scales. We extract the camera settings for $F_{r}$ frames using the CLIP text encoder, compute the differences, and then reshape the result into an embedding of size $F_{r}\times C\times H\times W$.
@@ 8.2 Embedding Encoder
9 More Details of Metrics
@@ 9.2 Consistency
9.3 Quality
We measure the quality of the generated frames by evaluating their alignment with the input prompts. Specifically, we use the CLIP [48] text and image encoders to obtain the features of the prompt and the generated frame, and then compute the cosine similarity between the two.
@@ 10 More Visual Results
We strongly encourage readers to view the videos (GIFs) on project page for more intuitive comparisons and additional case studies.

### v3

@@ 4.3 Ablation study
Differential Camera Encoder. To assess the impact of differential network design, we removed inter-frame differences in the differential camera encoder. Table 2 (w/o differential) suggests that incorporating different information enhances the model’s accuracy and consistency.
@@
Dataset Scaling. In Fig. 6, we present training dataset scaling experiments for shutter speed control. The results indicate that performance improves as the training data increases; however, beyond 1000 samples, the gains plateau. This suggests that our carefully designed dataset, combined with joint network optimization, is sufficient for understanding physical parameters and efficiently capturing fundamental principles with relatively little data.
Figure 6: Ablation study on dataset scaling characteristic.
@@ 5 Conclusions
Generative photography opens the door to a new paradigm in photography where the generated contents are consistent with the intrinsic camera physics. Our solution is based on two concepts: dimensionality lifting and differential camera intrinsics learning. It generates physically realistic camera effects while maintaining a high degree of scene consistency. Generative photography has the potential to reduce the post-processing burden in human-operated photography fundamentally. In addition, it offers a fresh perspective for generative models to understand the world better.
Acknowledgements. This work is supported, in part, by the National Science Foundation under grants 2133032, 2134209, 2030570, and 2431505.
Supplementary Material
@@ 6 Introduction
This supplementary material provides additional discussions and details on the construction of differential data (Section 7), network design (Section 8), evaluation metrics (Section 9), and more visual results (Section 10).
To better illustrate the continuity and effects of camera intrinsic setting control, we highly recommend readers view the Videos/GIFs provided in the project page: https://generative-photography.github.io/project/.
7 More Details of Building Differential Data
Our differential data pipeline dynamically generates training data by storing only base images and scene descriptions. Camera settings are sampled during training and simulated on-the-fly using physical principles, producing differential multi-frame data without pre-storing large video files. This also ensures continuous sampling of training data.
We provide below additional key considerations for constructing differential datasets for each type of camera setting, along with sample demonstrations.
7.1 Differential Data for Bokeh Rendering
@@
7.2 Differential Data for Focal Length
In the real world, obtaining a set of images of the same scene at multiple focal lengths is highly cumbersome, with a lack of perfect alignment between the images, and the achievable focal length range is limited [76]. In this paper, we reference the level-of-detail [64, 41] approach and compute the field-of-view (FoV) ratio of the desired focal length relative to the base image focal length. This ratio is then used for center cropping to approximate the actual continuous optical zoom process. In this subsection, we compare the performance of our method with that of actual optical zoom.
A camera’s field-of-view (FoV) can be expressed in terms of the focal length $f$ and the sensor dimensions (typically width $w$ or height $h$). The formulas are as follows:
Horizontal FoV:
@@ 7.2 Differential Data for Focal Length
7.3 Differential Data for Shutter Speed
@@
7.4 Differential Data for Color Temperature
@@
8 More Details of Differential Camera Encoder
In the Differential Camera Encoder, an important aspect is the incorporation of the differences in camera setting scales. We extract the camera settings for $F_{r}$ frames using the CLIP text encoder, compute the differences, and then reshape the result into an embedding of size $F_{r}\times C\times H\times W$.
@@ 8.2 Embedding Encoder
9 More Details of Proposed Metrics
@@ 9.2 Consistency
9.3 Following
We measure the prompt following of the generated frames by evaluating their alignment with the input prompts. Specifically, we use the CLIP [49] text and image encoders to obtain the features of the prompt and the generated frame, and then compute the cosine similarity between the two.
@@ 10 More Visual Results
We strongly encourage readers to view the videos/GIFs we provide for more intuitive comparisons and additional case studies.
@@
Figure 13: Visual comparisons between different generative methods on camera color temperature control. Both AnimateDiff [21] and CameraCtrl [23] have been fine-tuned/trained on our data.